# Beyond Training: Dynamic Token Merging for Zero-Shot Video Understanding

### v1

@@ 5 Analysis
5.1 Scaling Up Model Sizes
@@
As shown in Table 3, scaling up model sizes significantly enhances the performance of DyTo across structured VQA tasks, demonstrating notable gains over 7B performance shown in Table 1. Specifically, with a 34B model, DyTo achieves a 7.2% accuracy increase on NExTQA [32], an 8.2% boost on EgoSchema [24], and a 7.9% improvement on STAR [30]. These more substantial increments reflect the method’s enhanced capacity to capture and reason over complex spatial-temporal interactions, particularly in tasks requiring nuanced contextual comprehension.
@@
Compared to other training-free approaches at 34B, DyTo consistently outperforms IG-VLM [12] and SlowFast-LLaVA [35] across benchmarks, with accuracy advantages on tasks like VideoMME [8]. This suggests that the proposed hierarchical clustering and bipartite token merging mechanisms are particularly effective in leveraging the additional model capacity to retain critical semantic information, even in the absence of fine-tuning.
@@
In the open-ended VQA tasks shown in Table 4, our performance with a 34B model also surpasses its 7B counterpart in Table 2, underscoring the scalability of our approach. For instance, DyTo achieves a 2.3-point average increase across correctness and contextual understanding dimensions, as well as improvements in detail orientation and temporal understanding. These gains further highlight the robustness of DyTo adaptive framework, which scales efficiently with model size to provide more contextually enriched and temporally accurate responses across various VQA benchmarks.
Method LLM Video Vision NExTQA EgoSchema IntentQA STAR VideoMME
Size Trained Encoder acc. acc. acc. acc. acc.
VideoLLaMA2 46.7B CLIP-L O - 51.7 - - -
LLoVi GPT-3.5 Unknown X 67.7 50.3 64.0 - -
VideoAgent GPT-4 Unknown X 71.3 60.2 - - -
IG-VLM 34B CLIP-L X 70.9 53.6 65.3 51.0 52.0
SlowFast-LLaVA 34B CLIP-L X 71.9 55.8 66.2 50.7 53.2
DyTo 34B CLIP-L X 72.9 56.8 66.4 51.1 53.4
Table 3: Performance comparison on structured VQA Benchmarks with Larger Model Sizes (34B+).
Method LLM Vision Video MSVD-QA MSRVTT-QA TGIF-QA ANet-QA Video-ChatGPT (VCG Benchmark)
Size Encoder Trained acc./score acc./score acc./score acc./score CI DO CU TU CO Average
Video-LLAMA2 46.7B CLIP-L O 70.5/3.8 - - 50.3/3.4 3.08 3.11 3.64 2.67 3.26 3.15
PLLaVA 34B CLIP-L O 79.9/4.2 68.7/3.8 80.6/4.3 60.9/3.7 - - - - - -
LLaVA-NeXT-Video 34B CLIP-L O - - - 58.8/3.4 - - - - - -
LLaVA-NeXT-Image 34B CLIP-L X - - - 55.6/3.3 3.29 3.23 3.83 2.51 3.47 3.27
IG-VLM 34B CLIP-L X 79.6/4.1 62.4/3.5 79.1/4.2 58.4/3.5 3.21 2.87 3.54 2.51 3.34 3.09
SlowFast-LLaVA 34B CLIP-L X 78.7/4.1 67.1/3.7 80.6/4.3 58.8/3.5 3.48 2.96 3.84 2.70 3.54 3.30
DyTo-34B 34B CLIP-L X 79.6/4.1 66.1/3.6 80.7/4.2 59.0/3.5 3.45 2.95 3.82 2.72 3.63 3.32
Table 4: Performance comparison on open-ended VQA benchmarks with larger model sizes (34B+).
5.2 Performance on Various Video Lengths
@@
Figure 3: left: Performance comparison of baseline method under various video lengths. right: Effect of different input sampling lengths under various video lengths.
@@
We analyze the performance of DyTo across varying video lengths to better demonstrate its robustness, particularly in maintaining accuracy with longer videos. As shown in Fig. 3 left, while the performance of DyTo does show some degradation as video length increases, the decline is significantly less pronounced compared to other methods like IG-VLM [12] and SlowFast-LLaVA [35]. This stability is largely attributed to DyTo’s adaptive, video-dependent dynamic token merging, which effectively prioritizes essential frames and contextual tokens, preserving key information even as the video sequence lengthens.
In Fig. 3 right, we further examine the impact of increasing the number of sampled frames on performance with extended videos. As the number of sampled frames increases, DyTo’s accuracy on longer videos notably improves, demonstrating its capacity to capture detailed temporal and contextual information more effectively than competing methods. This improvement highlights the advantage of DyTo’s hierarchical clustering and bipartite token merging, which dynamically adjust to sample frames that represent critical moments, ensuring more comprehensive and accurate video understanding.
5.3 Visualizations of Hierarchical Clustering
5.4 Visualizations of Dynamic Bipartite Merging
@@ 5.5 Qualitative Case Study
Fig. 9 provides a qualitative illustration of DyTo’s robust performance in a complex zero-shot video understanding scenario from the NExTQA benchmark. In this instance, DyTo accurately identifies and preserves critical visual details, such as the red toy on the stroller, the child’s attire, and the architectural background. Due to the simple uniform sampling and pooling method, IG-VLM and SlowFast-LLaVA both output the hallucination content (carrying a backpack and holding a cup of cereal respectively). Unlike these methods, DyTo captures nuanced visual cues and contextual elements. DyTo’s hierarchical clustering, which segments video frames dynamically, alongside its bipartite merging strategy that avoids excessive compression that might lead to semantic loss, result in a more detailed and accurate scene understanding.
@@ Appendix A Time Consumption Experiment
Appendix B More Qualitative Cases
Figure 8: Example from STAR benchmark. red part denotes the incorrect or hallucination content. green part denotes the correct object content.

### v2

@@ 5 Analysis
5.1 Scalabity on Base Model Size
@@
As shown in Table 1, scaling up model sizes significantly enhances the performance of DyTo across structured VQA tasks, demonstrating notable gains over performance with 7B base models. Specifically, with a 34B model, DyTo achieves a 7.2% accuracy increase on NExTQA [36], an 8.2% boost on EgoSchema [28], and a 7.7% improvement on MVBench [16]. These more substantial increments reflect the method’s enhanced capacity to capture and reason over complex spatial-temporal interactions, particularly in tasks requiring nuanced contextual comprehension.
@@
Compared to other training-free approaches at 34B, DyTo consistently outperforms IG-VLM [11] and SlowFast-LLaVA [39] across benchmarks, with clear accuracy advantages on tasks like VideoMME [7]. This suggests that the proposed hierarchical clustering and bipartite token merging mechanisms are particularly effective in leveraging the additional model capacity to retain critical semantic information, even in the absence of fine-tuning.
@@
In the open-ended VQA tasks shown in Table 2, our performance with a 34B model also surpasses its 7B counterpart, reinforcing the scalability of our approach. For instance, DyTo achieves a 2.3-point average increase across correctness and contextual understanding dimensions of VCG benchmark, as well as improvements in detail orientation and temporal understanding. These gains further highlight the robustness of DyTo’s adaptive framework, which scales efficiently with model size to provide more contextually enriched and temporally accurate responses across various VQA benchmarks.
5.2 Performance w.r.t. Video Lengths
@@
Figure 3: Top: Performance comparison of baseline method under various video lengths. Bottom: Effect of different input sampling lengths under various video lengths.
We analyze the performance of DyTo across varying video lengths to better demonstrate its robustness, particularly in maintaining accuracy with longer videos. As shown in Fig. 3 (Top), while the performance of DyTo does show some degradation as video length increases, the decline is significantly less pronounced compared to other methods like IG-VLM [11] and SlowFast-LLaVA [39]. This stability is largely attributed to DyTo’s adaptive, video-dependent dynamic token merging, which effectively prioritizes essential frames and contextual tokens, preserving key information even as the video sequence lengthens.
In Fig. 3 (Bottom), we further examine the impact of increasing the number of sampled frames on performance with extended videos. As the number of sampled frames increases, DyTo’s accuracy on longer videos notably improves, demonstrating its capacity to capture detailed temporal and contextual information more effectively than competing methods. This improvement highlights the advantage of DyTo’s hierarchical clustering and bipartite token merging, which dynamically adjust to sample frames that represent critical moments, ensuring more comprehensive and accurate video understanding.
5.3 Ablation of Clustering and Merging Module
The experimental results from the table demonstrate significant differences in performance when DyTo adopts different strategies.verall, DyTo consistently outperforms both the baseline (SlowFast) and its ablated variants. Notably, on NExTQA, DyTo w/o Clustering slightly surpasses the full DyTo model. A closer inspection reveals that, due to the relatively short clips in NExTQA, the benefits of clustering become less pronounced, thus minimizing DyTo’s advantage. Nevertheless, in all other datasets, DyTo exhibits clear improvements over the baseline and ablation models, highlighting its stronger capability in handling more complex or longer video tasks. Furthermore, this performance edge persists regardless of the language model size (7B or 34B), demonstrating DyTo’s effectiveness and robustness in diverse settings.
5.4 Visualizations of Hierarchical Clustering
Method LLM NExTQA EgoSchema IntentQA VideoMME MVBench
Size acc. acc. acc. acc. acc.
SlowFast 7B 64.2 45.5 60.1 40.4 44.8
DyTo w/o Clustering 7B 65.6 47.8 61.4 42.3 44.8
DyTo w/o Token Merge 7B 64.9 45.6 60.5 41.2 44.6
DyTo 7B 65.7 48.6 61.6 42.7 45.2
SlowFast 34B 71.9 55.8 66.2 53.2 51.2
DyTo w/o Clustering 34B 73.2 55.8 66.8 52.5 52.7
DyTo w/o Token Merge 34B 72.2 56.0 66.3 51.7 51.8
DyTo 34B 72.9 56.8 67.3 53.4 52.9
Table 3: Ablation Study on clustering and token merging modules with parameter sizes ranging from 7B to 34B.
@@ 5.5 Qualitative Case Study
Fig. 9 provides a qualitative illustration of DyTo’s robust performance in a complex zero-shot video understanding scenario from the NExTQA benchmark. Based models is LLaVA-NeXT-Image-7B. In this instance, DyTo accurately identifies and preserves critical visual details, such as the red toy on the stroller, the child’s attire, and the architectural background. Due to the simple uniform sampling and pooling method, IG-VLM and SlowFast-LLaVA both output the hallucination content (carrying a backpack and holding a cup of cereal respectively). Unlike these methods, DyTo captures nuanced visual cues and contextual elements. DyTo’s hierarchical clustering, which segments video frames dynamically, alongside its bipartite merging strategy that avoids excessive compression that might lead to semantic loss, result in a more detailed and accurate scene understanding.
@@ Appendix A Time Consumption Experiment
Appendix B Implementation details
All evaluations could be conducted on a single Nvidia A100 80G graphics card. To accelerate inference, we use a Linux server equipped with 8 Nvidia A100 80G cards. We carry out our evaluation across three model series and five model size. The weights for these models are available on Huggingface [^1] https://huggingface.co/collections/liuhaotian/llava-16-65b9e40155f60fd046a5ccf2222https://huggingface.co/OpenGVLab/InternVL2333https://huggingface.co/Qwen/Qwen-VL-Chat. We implement rotary position embedding (RoPE) and apply a scaling factor of 2, extending the original context length from 4096 to 8192 tokens.
Appendix C Visualizations of Dynamic Bipartite Merging
Appendix D More Qualitative Cases
Figure 8: Example from MVBench benchmark. red part denotes the incorrect or hallucination content. green part denotes the correct object content.